# Optuna vs Code Llama: Are LLMs a New Paradigm for Hyperparameter Tuning?

### v1

@@ 6.2 Models from test dataset
Model Epochs RMSE Optuna RMSE FT 2 Difference
InceptionV3 1 0.4260 0.1741 0.2520
InceptionV3 2 0.5904 0.2339 0.3565
MaxVit 1 0.2120 0.3619 -0.1420
MaxVit 2 0.2306 0.4748 -0.2442
VisionTransformer 1 0.2673 0.4117 -0.1445
VisionTransformer 2 0.3045 0.4750 -0.1705
@@
For InceptionV3, fine-tuning in the second cycle yields substantial improvements. Both epochs exhibit significant reductions in RMSE compared to the Optuna baseline, with improvements of 0.2520 and 0.3565 for epochs 1 and 2, respectively. This highlights the effectiveness of fine-tuning in refining the hyperparameters for this model, allowing it to achieve a higher level of accuracy. Such results suggest that InceptionV3’s architecture is well-suited to benefit from the fine-tuning process, even when not included in the initial training dataset.
On the other hand, the performance of MaxVit and VisionTransformer models shows a different trend. For these models, fine-tuning leads to an increase in RMSE values for both epochs, with differences of -0.1420 and -0.2442 for MaxVit and -0.1445 and -0.1705 for VisionTransformer across the two epochs. This suggests that certain model architectures may not align as well with the specific hyperparameters derived from the fine-tuning process, potentially requiring further refinement or alternate optimization strategies.
@@ 7.1 Models from train dataset
The fine-tuning process demonstrates its ability to enhance accuracy for many models, with notable improvements observed during the second cycle. For instance, MNASNet and MobileNetV2 achieve significant gains in the first epoch of the second fine-tuning cycle, underscoring the effectiveness of fine-tuning in refining hyperparameters to better align with the models’ architectures. Similarly, GoogLeNet shows consistent accuracy improvements across cycles, particularly excelling in the second cycle, showcasing the iterative benefits of Code Llama’s fine-tuning framework.
@@
While certain models, such as MobileNetV3 and ConvNeXt, exhibit moderate or variable improvements across cycles, these findings underscore the importance of tailoring fine-tuning strategies to the specific requirements of each architecture. For models that are well-optimized in the initial cycle, additional fine-tuning may yield diminishing returns, suggesting the need for a balanced approach that considers computational costs alongside expected performance gains. This variability highlights the importance of model-specific strategies in leveraging the full potential of fine-tuning.
@@ 7.2 Models from test dataset
MaxVit demonstrates the most consistent and positive results among the evaluated models. The difference is minimal for one epoch (-0.0018), and for two epochs, FT Cycle 2 outperforms Optuna with a positive difference of +0.0249. This improvement underscores MaxVit’s adaptability to fine-tuning, highlighting its potential for further optimization. Such results position MaxVit as a strong candidate for deeper exploration in hyperparameter tuning and fine-tuning workflows.

### v2

@@ 6.2 Models from test dataset
Among the evaluated models, MaxVit and VisionTransformer show substantial improvements after fine-tuning. MaxVit achieves positive RMSE differences of 0.1919 and 0.2941 for epochs 1 and 2, respectively, while VisionTransformer improves by 0.1884 and 0.2406. These results indicate that the fine-tuning process significantly enhances the accuracy of these architectures, even though they were not part of the original training set. Such outcomes suggest that Code Llama generalizes well to new models, successfully adapting its hyperparameter recommendations to unfamiliar architectures.
In contrast, InceptionV3 demonstrates a notable decline in performance following the second fine-tuning cycle. RMSE values increase by –0.2363 and -0.3364 for epochs 1 and 2, respectively, compared to the Optuna baseline. While these results indicate that the current fine-tuning setup may not fully optimize InceptionV3, they also point to valuable opportunities for future enhancement. With refined prompt design, adjusted parameter initialization, or extended fine-tuning schedules, there is strong potential to better align the model’s architecture with the optimization process and achieve improved outcomes.
@@ 7.1 Models from train dataset
The fine-tuning process demonstrates its ability to enhance accuracy for many models, with notable improvements observed during the second cycle. For instance, MNASNet and MobileNetV2 achieve significant gains in the first epoch of the second fine-tuning cycle, underscoring the effectiveness of fine-tuning in refining hyperparameters to better align with the models’ architectures. Similarly, RegNet and SqueezeNet show consistent accuracy improvements across cycles, particularly excelling in the second cycle, showcasing the iterative benefits of Code Llama’s fine-tuning framework.
@@
While certain models, such as MobileNetV3 and SwinTransformer, exhibit moderate or variable improvements across cycles, these findings underscore the importance of tailoring fine-tuning strategies to the specific requirements of each architecture. For models that are well-optimized in the initial cycle, additional fine-tuning may yield diminishing returns, suggesting the need for a balanced approach that considers computational costs alongside expected performance gains. This variability highlights the importance of model-specific strategies in leveraging the full potential of fine-tuning.
@@ 7.2 Models from test dataset
MaxVit demonstrates the most consistent and positive results among the evaluated models. The difference is minimal for one epoch (-0.0018), and for two epochs, FT Cycle 2 outperforms Optuna with a positive difference of 0.0249. This improvement underscores MaxVit’s adaptability to fine-tuning, highlighting its potential for further optimization.